# Evaluating the Robustness of Adversarial Defenses in Malware Detection Systems

### v1

@@ II-C Evasion Attacks
Adversarial attacks targeting malware detection systems can be broadly classified into oblivious and adaptive attacks, based on the knowledge the attacker has of the adversary detector $g$.
@@
Oblivious attacks operate without any knowledge of the adversary detector $g$. These attacks aim solely to bypass the malware detector $f$ by ensuring that adversarial inputs are misclassified as benign. As such, oblivious attacks overlook the presence of $g$, making them less effective when $g$ is actively present. The adversary detector $g$ can flag such inputs as adversarial, limiting their ability to evade the overall detection framework.
@@
Adaptive attacks, on the other hand, explicitly account for the adversary detector $g$, represented as $g(z^{\prime})=\psi_{\vartheta}(\phi(z^{\prime}))$. These attacks require that the adversarial input not only evades the malware detector $f$ but also satisfies $g(z^{\prime})\leq\tau$. This dual requirement makes adaptive attacks more complex and challenging to defend against.
The optimization problem for adaptive attacks is formulated  as:
@@
This formulation substitutes the condition $\varphi_{\theta}(x^{\prime})=0$ with minimizing $\mathcal{L}_{\varphi}(\theta,x^{\prime},0)$, addressing the issue of non-differentiability.
A critical challenge in this optimization process lies in the non-linear nature of the constraint $\psi_{\vartheta}(x^{\prime})\leq\tau$, which precludes the use of standard gradient descent techniques. Unlike oblivious attacks, the problem cannot be reformulated by simply swapping the objective and constraint. The non-linear constraint imposed by $g$ requires a more sophisticated approach to ensure that adversarial inputs evade both $f$ and $g$ while adhering to distortion bounds.
To address this difficulty, existing literature adopts a Lagrangian relaxation approach, previously applied to constructing minimum-distortion adversarial examples [28]. The optimization objective is reformulated as:
@@
where $\lambda\geq 0$ is a penalty factor that adjusts the relative importance of evading $f$ versus reducing the response of $g$. The optimal value of $\lambda$ is determined through a binary search, ensuring an effective trade-off between these competing objectives [29].
@@
II-E Binary Rounding Methods
@@
  • Deterministic Binary Rounding: This method applies a fixed threshold, typically 0.5, to convert continuous values into binary features. Values equal to or greater than 0.5 are rounded to 1, while values below 0.5 are rounded to 0.
@@ III Proposed Approach
III-A Threat Model and Design Objective
We assume a white-box attack setting, where the attacker has full knowledge of both the malware detector $f$ and the adversary detector $g$, including their architectures, trained parameters, and feature representations [28, 30]. This assumption allows for worst-case evaluations of the robustness of ML-based defenses against adversarial attacks.
The attacker’s objective is to generate adversarial examples that evade $f$ while remaining undetected by $g$. An approximate inverse mapping function $\tilde{\phi}^{-1}$ is employed, as proposed in prior studies [23, 20], to translate feature-space perturbations into the problem space, ensuring the adversarial examples retain their functional validity.
@@ III-A Threat Model and Design Objective
III-B Incorporating Distance Metrics
@@
III-C Problem Formulation
The adversarial attack is formulated as an optimization problem in the binary feature space. For a given malware instance-label pair $(x,y)$, where $x=\phi(z)$ and $y=1$, the attacker seeks to determine the minimal feature modification $\delta^{\star}$, quantified by the Hamming distance $d_{H}$, such that the resulting adversarial example $x^{\star}=x+\delta^{\star}$ satisfies:
@@ III-C Problem Formulation
$\quad x+\delta\in[\underline{u},\overline{u}],$ (13)
@@
[…]
Where $\widetilde{d}_{H}(x,x+\delta)$ represents a differentiable approximation of the Hamming distance, normalized by the number of features $d$ to ensure its value lies within the interval $[0,1]$. This normalization removes the need for hyperparameter tuning to balance the trade-off between the loss and the perturbation size, thereby avoiding computationally expensive line searches for each input sample [31].
The total loss $\mathcal{L}$ is defined as:
@@
  • $\kappa_{1}$ and $\kappa_{2}$ are confidence margins for $f$ and $g$, ensuring robust misclassification and detection avoidance.
We define the loss for the malware detector $f$ using the differences between logits, a method whose rationale is extensively discussed in [28]. Furthermore, independently bounding the loss functions ensures that minimizing one loss term does not inherently satisfy the constraints for both detectors, thereby preserving the distinct objectives of $f$ and $g$.
III-C1 Hamming Distance Approximation
To enable gradient-based optimization, the Hamming distance $d_{H}$ is approximated as:
@@ III-C1 Hamming Distance Approximation
Here, $\sigma>0$ serves as a smoothing parameter to ensure differentiability. However, this approximation may lead to non-sparse solutions. To mitigate this, an adaptive projection operator $\Pi_{\gamma}$ enforces sparsity by setting perturbation components below a threshold $\gamma$ to zero, promoting both sparsity and adversarial effectiveness [31].
@@
Loss$(\theta,\vartheta,x,\delta)=\mathcal{L}(\theta,\vartheta,x+\delta)+\frac{1}{d}\widetilde{d}_{H}(x,x+\delta)$;
@@
             if $(x+\delta)$ evades both $f$ and $g$ then
@@
[…]
             $(x+\delta_{\text{binary}})$ evades both $f$ $g$ then
@@
[…]
                   if $(x+\delta)$ evades both $f$ $g$ then
@@
III-C2 Binary Rounding
To convert continuous solutions into binary values, we employ Prioritized Binary Rounding (detailed in Subsection III-D), which prioritizes feature modifications based on their impact on model decisions. This method ensures compliance with binary constraints while minimizing perturbations.
III-C3 Dynamic Confidence Adjustment
In adversarial optimization, the solution typically converges to an optimal result in the continuous-valued space. However, during the binary rounding process, where continuous values are discretized into binary form, suboptimality can be introduced. This suboptimality arises because the process of forcing discrete values can disrupt the fine-tuned structure of the continuous solution. To address this issue, we employ the Dynamic Confidence Adjustment method. After the initial convergence of the optimization process, the confidence parameter $\kappa$ is dynamically adjusted within a bounded range $[-\alpha,+\alpha]$. This adjustment enables the optimization process to explore regions near the decision boundary, facilitating the discovery of a binary solution that more effectively satisfies the adversarial objectives.
III-C4 Algorithm Details for the $\sigma$-binary Attack
@@
The algorithm employs two nested loops: an outer binary search loop to adjust $C$ for balancing loss terms, and an inner optimization loop to minimize a loss function combining the evasion objective and a smooth approximation of the Hamming distance. Perturbations are iteratively updated via a cosine annealing learning rate, clipped gradient descent, and enforced sparsity. Non-binary perturbations are converted to binary form using prioritized binary rounding to maintain functionality and input constraints.
@@
III-D Prioritized Binary Rounding
To efficiently map continuous perturbations to binary feature vectors, the proposed $\sigma$-binary attack framework introduces a novel technique called Prioritized Binary Rounding. This method ensures compliance with binary constraints while minimizing unnecessary feature modifications, balancing adversarial effectiveness and sparsity.
Traditional binary rounding methods often use uniform deterministic or randomized thresholds, which overlook variations in feature importance and adversarial impact. In contrast, Prioritized Binary Rounding selects features for modification based on two critical factors:
  • Perturbation Magnitude: Features exhibiting larger deviations from their original values are prioritized for modification.
  • Feature Importance: Features that significantly influence the model’s decision boundary are prioritized to enhance adversarial effectiveness.
@@ III-D Prioritized Binary Rounding
Compute perturbation magnitude: $\mathbf{\Delta}\leftarrow|\mathbf{x}_{\text{adv}}-\mathbf{x}_{\text{orig}}|$;
@@
By leveraging this method, the $\sigma$-binary attack achieves higher success rates with minimal perturbations, making it a robust and efficient solution for adversarial attacks in malware detection scenarios.
IV Experiments
To evaluate the effectiveness of our proposed methods, we conducted extensive experiments across various scenarios. Our analysis addresses two primary categories of research questions (RQs): (i) comparing the proposed methods with prior work (RQ1 and RQ2), and (ii) using these methods to evaluate the performance of existing defenses (RQ3, RQ4, and RQ5). Specifically, we sought answers to the following:
  • RQ1: Effectiveness of Prioritized Binary Rounding. How effective is the proposed Prioritized Binary Rounding technique in comparison to existing binary rounding methods?
@@ IV Experiments
  • RQ2: Effectiveness of the $\sigma$-binary attack. How does our proposed $\sigma$-binary attack perform relative to other attack strategies in binary feature spaces?
  • RQ3: Baseline performance of defenses. How effective are defense mechanisms under non-adversarial conditions?
@@
  • RQ4: Robustness against oblivious attacks. How resilient are defenses against oblivious attacks, where attackers are unaware of the adversary detector $g$?
  • RQ5: Robustness against adaptive attacks. How robust are the defenses against adaptive attacks? This evaluates performance under worst-case scenarios.
IV-A Experimental Framework and Configuration
This section provides a detailed overview of the dataset, feature extraction process, defense mechanisms, attack methods, evaluation metrics, and experimental configurations utilized in our study.
IV-A1 Dataset
@@
We utilized the MalScan dataset [26], a widely recognized collection comprising 23,196 Android applications, with 11,583 malicious and 11,613 benign samples. The dataset spans applications from 2011 to 2018. The data was randomly split into training (60%), validation (20%), and testing (20%) subsets.
IV-A2 Feature Extraction
@@
Feature extraction was performed using Drebin [4], which analyzes Android application packages (APKs) to construct a binary feature space. Features were extracted from the Android manifest and the disassembled dexcode using the Androguard tool. These features were organized into eight categories, including hardware components, permissions, intents, API calls, and class names. Each APK is represented as a binary feature vector, where each dimension indicates the presence ($1$) or absence (00) of a specific feature. Consistent with prior work [23], we exclude easily modifiable features (e.g., package names) and retain the 10,000 most frequent features, ensuring robust representation for evaluation.
IV-A3 Defenses Considered for Comparative Analysis
@@
Below, we provide a detailed overview of each defense:
@@
  • DNN [8]: A baseline deep neural network model for malware detection. This model does not incorporate specific countermeasures against evasion attacks and serves as a reference for comparing the robustness of advanced defense mechanisms.
@@
  • DNN+ [32]: An extension of the DNN model that incorporates a secondary detector. This detector introduces an additional outlier class to identify adversarial examples.
@@
  • KDE [33]: Combines a DNN with a Kernel Density Estimation (KDE)-based secondary detector. The KDE mechanism identifies adversarial examples by analyzing deviations in activations of the DNN’s penultimate layer.
@@
  • DLA [34]: Dense Layer Analysis (DLA) incorporates a secondary detector alongside the DNN. By evaluating activations across all dense layers, DLA distinguishes between normal and adversarial examples, adding a robust layer of defense.
@@
  • AT-rFGSMk [9]: A defense mechanism that strengthens the DNN through adversarial training using the FGSMk attack. This approach employs randomized rounding projections to increase the model’s resilience to adversarial perturbations.
@@
  • AT-MaxMA [20]: Enhances DNN robustness by incorporating adversarial training with a Mixture of Attacks (MaxMA) strategy. This method combines multiple attack techniques to construct a more comprehensive and resilient defense.
@@
  • ICNN [23]: Integrates a DNN with an Input Convex Neural Network (ICNN) as a secondary detector. The ICNN operates independently of the DNN’s architecture and analyzes the feature space to detect adversarial examples.
@@
  • PAD-SMA [23]: integrates a DNN-based malware detector with an ICNN-based adversary detector. Both components are fortified through adversarial training using the Stepwise Mixture of Attacks (SMA), making this approach a representative example of defenses that combine adversarial training with adversary detection mechanisms.
This comprehensive classification facilitates a thorough analysis of the robustness of various defenses against the proposed $\sigma$-binary attack.
IV-A4 Evasion Attack Methods for Comparative Analysis
To benchmark our proposed $\sigma$-binary attack, we compare it against four well-established adversarial attack methods: CW, $\sigma$-zero, Mimicry, and PGD. Below, we detail each method:
Projected Gradient Descent (PGD) Attacks
PGD [35] is an iterative optimization method that identifies perturbations maximizing adversarial loss while adhering to predefined constraints. The update rule is expressed as:
$\delta^{(t+1)}_{x}=\text{Proj}_{[u-x,u-x]}\Big{(}\delta^{(t)}_{x}+\alpha\nabla_{\delta_{x}}F(\theta,x+\delta^{(t)}_{x},1)\Big{)},$
where $\alpha>0$ is the step size, $\nabla_{\delta_{x}}F$ denotes the gradient of the loss function, and Proj ensures that perturbations remain within constraint bounds. To address small gradients and enhance adversarial success, researchers employ normalization in different norms, such as the $\ell_{\infty}$-norm, $\ell_{2}$-norm, or $\ell_{1}$-norm. Each norm identifies the steepest direction, which is then utilized to compute effective perturbations  [36].
CW Attack
The Carlini & Wagner (CW) attack [28] minimizes the $\ell_{2}$-norm of perturbations while ensuring misclassification. It achieves this by solving a carefully crafted optimization problem through gradient-based methods, resulting in highly effective evasion with minimal feature modification.
$\sigma$-zero Attack
The $\sigma$-zero attack [31] is designed to minimize the number of modified features by approximating the $\ell_{0}$-norm and applying adaptive projections. It is effective in generating highly sparse perturbations, challenging defenses that assume adversarial examples require extensive feature modifications.
Mimicry Attack
Mimicry [37, 27] is a gradient-free adversarial attack that perturbs malware samples to resemble benign samples. By iteratively querying the target model and adjusting perturbations, it effectively bypasses detection mechanisms without requiring knowledge of model parameters.
IV-A5 Evaluation Metrics

### v2

@@ III-D Adversarial Attack Methods
Adversarial attacks targeting malware detection systems can be broadly classified into oblivious and adaptive attacks, based on the attacker’s knowledge of the adversary detector $g$.
@@
Oblivious attacks operate without the knowledge of the adversary detector $g$. These attacks aim solely to bypass the malware detector $f$ by ensuring that adversarial inputs are misclassified as benign. As such, oblivious attacks overlook the presence of $g$, making them less effective when $g$ is actively present. The adversary detector $g$ can flag such inputs as adversarial, limiting their ability to evade the overall detection framework.
@@
Adaptive attacks, on the other hand, explicitly account for the adversary detector $g$, represented as $g(z^{\prime})=\psi_{\vartheta}(\phi(z^{\prime}))$. These attacks require that adversarial input not only evades the malware detector $f$ but also satisfies $g(z^{\prime})\leq\tau$. This dual requirement makes adaptive attacks more complex and challenging to defend against.
The optimization problem for adaptive attacks is formulated as follows:
@@
This formulation substitutes the condition $\varphi_{\theta}(x^{\prime})=0$ with the minimization of $\mathcal{L}_{\varphi}(\theta,x^{\prime},0)$, addressing the issue of non-differentiability.
A critical challenge in this optimization process lies in the nonlinear nature of the constraint $\psi_{\vartheta}(x^{\prime})\leq\tau$, which precludes the use of standard gradient descent techniques. Unlike oblivious attacks, the problem cannot be reformulated by simply swapping the objective and constraint. The non-linear constraint imposed by $g$ requires a more sophisticated approach to ensure that adversarial inputs evade both $f$ and $g$ while adhering to distortion bounds.
To address this difficulty, existing literature adopts a Lagrangian relaxation approach, previously applied to constructing minimum-distortion adversarial examples [carlini2017towards]. The optimization objective is reformulated as follows.
@@
where $\lambda\geq 0$ is a penalty factor that adjusts the relative importance of evading $f$ versus reducing the response of $g$. The optimal value of $\lambda$ is determined through a binary search, ensuring an effective trade-off between these competing objectives [bryniarski2021evading].
@@
III-F Binary Rounding Methods
@@
  • Deterministic Binary Rounding: This method applies a fixed threshold—typically 0.5—to convert continuous values into binary features. Values equal to or greater than 0.5 are rounded to 1, while values below 0.5 are rounded to 0.
@@ IV Proposed Approach
IV-A Threat Model and Design Objective
We assume a white-box attack setting, where the attacker has full knowledge of both the malware detector $f$ and the adversary detector $g$, including their architectures, trained parameters, and feature representations [carlini2017towards, biggio2018wild]. This assumption allows for worst-case evaluations of the robustness of ML-based defenses against adversarial attacks.
The attacker’s objective is to generate adversarial examples that evade $f$ while remaining undetected by $g$. An approximate inverse mapping function $\tilde{\phi}^{-1}$ is employed, as proposed in prior studies [li2023pad, li2020adversarial], to translate feature-space perturbations into the problem space, ensuring the adversarial examples retain their functional validity.
@@ IV-A Threat Model and Design Objective
IV-B Incorporating Distance Metrics
@@
IV-C Problem Formulation
The adversarial attack is formulated as an optimization problem in the binary feature space. For a given malware instance-label pair $(x,y)$, where $x=\phi(z)$ and $y=1$, the attacker seeks to determine the minimal feature modification $\delta^{\star}$, quantified by the Hamming distance $d_{H}$, such that the resulting adversarial example $x^{\star}=x+\delta^{\star}$ satisfies.
@@ IV-C Problem Formulation
$\quad x+\delta\in[\underline{u},\overline{u}].$ (13)
@@
$\quad\text{s.t.}\quad x+\delta\in\mathcal{X},\quad x+\delta\in[\underline{u},\overline{u}],$ (15)
Where $\widetilde{d}_{H}(x,x+\delta)$ represents a differentiable approximation of the Hamming distance, normalized by the number of features $d$ to ensure its value lies within the interval $[0,1]$. This normalization removes the need for hyperparameter tuning to balance the trade-off between the loss and the perturbation size, thereby avoiding computationally expensive line searches for each input sample [cina2024sigma].
The total loss $\mathcal{L}$ is defined as follows.
@@
  • $\kappa_{1}$ and $\kappa_{2}$ are confidence margins for $f$ and $g$, ensuring robust misclassification and avoidance of detection.
We define the loss for the malware detector $f$ using the differences between logits, a method whose rationale is extensively discussed in [carlini2017towards]. Furthermore, independently bounding the loss functions ensures that minimizing one loss term does not inherently satisfy the constraints for both detectors, thereby preserving the distinct objectives of $f$ and $g$.
IV-C1 Hamming Distance Approximation
To enable gradient-based optimization, the Hamming distance $d_{H}$ is approximated as follows.
@@ IV-C1 Hamming Distance Approximation
Here, $\sigma>0$ serves as a smoothing parameter to ensure differentiability. However, this approximation may lead to non-sparse solutions. To mitigate this, an adaptive projection operator $\Pi_{\gamma}$ enforces sparsity by setting perturbation components below a threshold $\gamma$ to zero, promoting both sparsity and adversarial effectiveness [cina2024sigma].
@@
Initialize: $\delta^{\star}\leftarrow\infty$, $C\leftarrow C_{\text{init}}$, best_dist $\leftarrow\infty$; for $\text{OuterStep}\leftarrow 1$ to $N_{\text{binary\_search}}$ do
@@
    if $(x+\delta)$ evades both $f$ & $g$ then
@@
     if $d_{H}(\delta_{\text{binary}},0)<\texttt{\small best\_dist}$ and
     $(x+\delta_{\text{binary}})$ evades both $f$ & $g$ then
@@
    if $|\texttt{\small Loss}-\texttt{\small previous\_Loss}|<\epsilon$ then
       if $(x+\delta)$ evades both $f$ & $g$ then
@@
IV-C2 Binary Rounding
To convert continuous solutions into binary values, we employ Prioritized Binary Rounding (detailed in Subsection IV-D), which prioritizes feature modifications based on their impact on model decisions. This method ensures compliance with binary constraints while minimizing perturbations.
IV-C3 Dynamic Confidence Adjustment
In adversarial optimization, the solution typically converges to an optimal result in the continuous-valued space. However, during the binary rounding process, where continuous values are discretized into binary form, suboptimality can be introduced. This suboptimality arises because the process of forcing discrete values can disrupt the fine-tuned structure of the continuous solution. To address this issue, we employ the Dynamic Confidence Adjustment method. After the initial convergence of the optimization process, the confidence parameter $\kappa$ is dynamically adjusted within a bounded range $\alpha,\alpha]$. This adjustment enables the optimization process to explore regions near the decision boundary, facilitating the discovery of a binary solution that more effectively satisfies the adversarial objectives.
IV-C4 Algorithm Details for the $\sigma$-binary Attack
@@
The algorithm employs two nested loops: an outer binary search loop to adjust $C$ for balancing loss terms, and an inner optimization loop to minimize a loss function combining the evasion objective and a smooth approximation of the Hamming distance. Perturbations are iteratively updated via a cosine annealing learning rate, clipped gradient descent, and enforced sparsity. Non-binary perturbations are converted to binary form using PBR to maintain functionality and input constraints.
@@
IV-D Prioritized Binary Rounding
To efficiently map continuous perturbations to binary feature vectors, the proposed $\sigma$-binary attack framework introduces a novel technique termed PBR. This method enforces binary constraints while minimizing unnecessary feature modifications, achieving a balance between adversarial effectiveness and perturbation sparsity.
Conventional binary rounding approaches typically rely on uniform deterministic or randomized thresholds, thereby neglecting variations in feature importance and their respective impact on the model’s decision. In contrast, PBR selects features for modification according to a principled ordering that combines two criteria:
  • Perturbation magnitude: Features exhibiting larger deviations from their original values are given higher priority for modification.
  • Feature importance: Features with greater influence on the model’s decision boundary are prioritized to strengthen the adversarial effect.
In the proposed implementation, features are primarily ranked by their gradient-derived perturbation magnitude. When multiple features exhibit similar magnitudes, discriminative feature importance—computed on the training dataset—is used as a secondary tie-breaking criterion. This auxiliary ranking improves rounding stability by prioritizing features that are statistically more informative for classification. Discriminative importance is estimated from the frequency of feature occurrence across classes, where features appearing predominantly in either benign or malicious samples are considered more discriminative. Consequently, the PBR mechanism prioritizes features that exhibit strong gradient magnitudes and high discriminative relevance to classification outcomes. The algorithm proceeds as follows.
@@ IV-D Prioritized Binary Rounding
Compute perturbation magnitude: $\boldsymbol{\Delta}\leftarrow|\mathbf{x}_{\text{adv}}-\mathbf{x}_{\text{orig}}|$;
@@
The joint use of $\sigma$-binary and PBR enables both fine-grained optimization and faithful binary constraint enforcement. This dual mechanism forms the foundation for the empirical evaluations presented in Section V.
V Experiments
We evaluate the proposed methods through an extensive experimental campaign designed to: (i) compare our adversarial attack against prior approaches, (ii) assess the robustness of multiple defense mechanisms under diverse threat models, and (iii) quantify trade-offs between robustness, standard classification performance, and computational cost. The evaluation is organized around five research questions (RQs):
@@ V Experiments
  • RQ1 — Performance of the $\sigma$-binary attack. How does the $\sigma$-binary attack—enhanced through the PBR mechanism—perform compared to existing adversarial attack strategies in binary feature spaces?
  • RQ2 — Baseline performance of defenses. What are the detection accuracies and false alarm characteristics of the evaluated defenses under standard (non-adversarial) conditions?
@@
  • RQ3 — Robustness to oblivious attacks. How resilient are defenses when the adversary is unaware of the adversary detector $g$ (i.e., the oblivious threat model)?
  • RQ4 — Robustness to adaptive attacks. How do defenses withstand adaptive attacks that explicitly account for both the malware detector $f$ and the adversary detector $g$, representing a worst-case threat model?
  • RQ5 — Trade-offs: robustness, accuracy, and computational cost. What are the empirical trade-offs between adversarial robustness, standard (non-adversarial) classification accuracy, and the computational cost of each defense strategy?
V-A Experimental framework
This subsection summarizes the dataset, feature extraction pipeline, defenses, and attacks evaluated, metrics, and experimental configuration used throughout the study.
V-A1 Dataset
@@
We utilized the MalScan dataset [wu2019malscan], a widely recognized collection comprising 30,715 Android applications, with 15,430 malicious and 15,285 benign samples. The dataset spans applications from 2011 to 2018 (Table I). The near-balanced class distribution mitigates training bias, while the eight-year span of samples captures temporal variability in benign and malicious behaviors.
Table I: Dataset summary (MalScan).
Year Benign Malware Total
2011 1,920 1,916 3,836
2012 1,875 2,000 3,875
2013 1,896 2,000 3,896
2014 1,826 1,982 3,808
2015 1,811 1,839 3,650
2016 2,015 1,940 3,955
2017 1,884 1,834 3,718
2018 2,058 1,919 3,977
Total 15,285 15,430 30,715
V-A2 Feature Extraction
@@
Feature extraction was performed using Drebin [arp2014drebin], which analyzes Android application packages (APKs) to construct a binary feature space. Features were extracted from the Android manifest and the disassembled DEX code using the Androguard tool. These features were organized into eight categories, such as hardware components, permissions, intents, API calls, and class names. Each APK is represented as a binary feature vector, where each dimension indicates the presence ($1$) or absence (0) of a specific feature. Consistent with prior work [li2023pad], we exclude easily modifiable features (e.g., package names) and retain the 10,000 most frequent features, ensuring robust representation for evaluation.
V-A3 Defenses Considered for Comparative Analysis
@@
Below, we provide a detailed overview of each defense.
@@
  • DNN [grosse2017adversarial]: A baseline deep neural network model for malware detection. This model does not incorporate specific countermeasures against evasion attacks and serves as a reference for comparing the robustness of advanced defense mechanisms.
@@
  • DNN+ [grosse2017statistical]: An extension of the DNN model that incorporates a secondary detector. This detector introduces an additional outlier class to identify adversarial examples.
@@
  • KDE [pang2018towards]: Combines a DNN with a Kernel Density Estimation (KDE)-based secondary detector. The KDE mechanism identifies adversarial examples by analyzing deviations in activations of the DNN’s penultimate layer.
@@
  • DLA [sperl2020dla]: Dense Layer Analysis (DLA) incorporates a secondary detector alongside the DNN. By evaluating activations across all dense layers, DLA distinguishes between normal and adversarial examples, adding a robust layer of defense.
@@
  • AT-rFGSMk [huang2018adversarial]: A defense mechanism that strengthens the DNN through adversarial training using the FGSMk attack. This approach employs randomized rounding projections to increase the model’s resilience to adversarial perturbations.
@@
  • AT-MaxMA [li2020adversarial]: Enhances DNN robustness by incorporating adversarial training with a Mixture of Attacks (MaxMA) strategy. This method combines multiple attack techniques to construct a more comprehensive and resilient defense.
@@
  • ICNN [li2023pad]: Integrates a DNN with an Input Convex Neural Network (ICNN) as a secondary detector. The ICNN operates independently of the DNN’s architecture and analyzes the feature-space to detect adversarial examples.
@@
  • PAD-SMA [li2023pad]: The PAD framework (Principled Adversarial Malware Detection) integrates a DNN-based malware detector with an ICNN-based adversary detector. Both components are fortified through adversarial training using the Stepwise Mixture of Attacks (SMA), making this approach a representative example of defenses that combine adversarial training with adversary detection mechanisms.
V-A4 Evasion attack methods for comparative analysis
To assess the effectiveness of the proposed $\sigma$-binary attack, we compare it against a representative and diverse suite of established evasion techniques encompassing gradient-based, ensemble, gradient-free, and optimization-driven approaches. The selection follows established benchmarking practice [li2023pad, li2020adversarial] and augmented with sparsity-oriented solvers for a more comprehensive analysis. Details on implementation are provided in Section III-D. The comparative set comprises:
  • Gradient-based methods: discrete-tailored and continuous gradient approaches, including Bit Coordinate Ascent (BCA), Bit Gradient Ascent (BGA), the Grosse gradient-directed addition, randomized FGSM (rFGSM), and PGD variants ($\ell_{1}, PGD$-$\ell_{2}$, PGD-$\ell_{\infty}$). Continuous outputs are converted to binary vectors via deterministic or randomized rounding as appropriate.
  • Ensemble / mixture attacks: Iterative MaxMA (iMaxMA) and Stepwise Mixture of Attacks (SMA), which combine multiple primitives to improve transferability [li2020adversarial, li2023pad].
  • Gradient-free methods: Mimicry-style attacks that modify malware features to resemble benign instances without requiring gradient access [vsrndic2014practical, biggio2013evasion].
  • Optimization / sparsity-oriented: Carlini–Wagner (CW, $\ell_{2}$) and the $\sigma$-zero solver targeting $\ell_{0}$-like sparsity [carlini2017towards, cina2024sigma].
Including both CW and $\sigma$-zero permits a direct comparison between our binary-targeted $\sigma$-binary attack and established distance-optimized methods that explicitly minimize perturbation magnitude.
V-A5 Evaluation Metrics